# Michael–Simon inequality for anisotropic energies close to the area via multilinear Kakeya-type bounds

### v1

@@ 3.3. Complementary inequalities
The vector field $S_{f}=\alpha_{f}Z$ admits a useful estimate on the complement $\mathcal{U}_{j}^{x}\setminus G_{j}^{x}$, as we now show.
@@
$\int_{\mathcal{U}_{j}^{x}\setminus G_{j}^{x}}(S_{f}^{x}-|S_{f}^{y}|)^{+}\leq\int_{\mathcal{U}_{j}^{x}}(S_{f}^{x}-|S_{f}^{y}|)^{-},$
as well as the analogous one for $T_{f}$ on $\mathcal{U}_{k}^{y}$ (with $x$ and $y$ interchanged).
This fact will be a direct consequence of the next elementary lemma.
@@
Assume now $\xi(L)>\xi(0)$. The image $\xi(E)$ includes $\xi(0),\xi(L)]$ since, for any $\xi(0)\leq\lambda\leq\xi(L)$, we have $\max f^{-1}(\lambda)\in E$. Since $E\subseteq\{\dot{\xi}\geq 0\}$ up to negligible sets, the area formula gives
@@
As above, we write $S_{f}=\nabla^{\perp}\varphi$. Given a regular level set of $\varphi$ (where $\alpha_{f}>0$), we can parametrize it with an integral curve $\gamma:[0,L]\to\mathcal{U}_{j}^{x}$ of $Z/|Z|$, with $\gamma^{x}(0)=j$ and $\gamma^{x}(L)=j+1$. Note that $\gamma$ has unit speed and that
@@
We apply the previous lemma with $\xi(t):=\int_{0}^{t}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)$. With $E$ as in the statement of the lemma, if $s\in E$ then for all $t>s$ we have
@@
Integrating over all regular level sets of $\varphi$ and using the coarea formula, we get
$\int_{\mathcal{U}_{j}^{x}\setminus G_{j}^{x}}(-\partial_{y}\varphi-|\partial_{x}\varphi|)^{+}\leq\int_{\mathcal{U}_{j}^{x}}(-\partial_{y}\varphi-|\partial_{x}\varphi|)^{-}.$
@@
Recall that $0\leq\alpha_{f}\leq\alpha$, which gives $0\leq\alpha_{d}\leq\alpha$.
@@
For any $\Lambda>2$, we have
$\int_{\mathcal{U}_{j}^{x}}(S_{d}^{x}-|S_{d}^{y}|)^{+}\leq\Lambda\int_{\mathcal{U}_{j}^{x}}\lvert\operatorname{div}S\rvert+\frac{1}{\Lambda-2}\int_{\mathcal{U}_{j}^{x}}(S_{d}^{x}-|S_{d}^{y}|)^{-}.$
@@
Writing $Z=\nabla^{\perp}f$, consider a level set $\gamma:[0,L]\to\mathcal{U}_{j}^{x}$ of $f$, parametrized by arclength. We have
@@
where we write $\alpha_{d}$ in place of $\alpha_{d}\circ\gamma$ for simplicity. We distinguish two cases: if the length $\mathcal{H}^{1}(\gamma\cap\{\alpha_{d}>s\})\leq\Lambda$, we just bound the inner integral by $\Lambda$.
Otherwise, note that, since $\dot{\gamma}^{x}+|\dot{\gamma}^{y}|\geq 1$ and $\int_{\{\alpha_{d}>s\}}\dot{\gamma}^{x}\leq\int_{0}^{L}\dot{\gamma}^{x}=1$ (as $\dot{\gamma}^{x}>0$), we must have
$\int_{\{\alpha_{d}>s\}}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\geq\int_{\{\alpha_{d}>s\}}(|\dot{\gamma}^{y}|-\dot{\gamma}^{x})\geq\int_{\{\alpha_{d}>s\}}(1-2\dot{\gamma}^{x})\geq\Lambda-2,$
which gives
$\int_{\{\alpha_{d}>s\}}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\leq\int_{\{\alpha_{d}>s\}}\dot{\gamma}^{x}\leq 1\leq\frac{1}{\Lambda-2}\int_{\{\alpha_{d}>s\}}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}.$
Also, the maximum of $\alpha_{d}$ along $\gamma$ is bounded by $\int_{0}^{L}|\dot{\alpha}_{d}|$. Since $\dot{\gamma}=\frac{Z}{|Z|}$ and $\operatorname{div}Z=0$, we have $|\dot{\alpha}_{d}|=\frac{\lvert\operatorname{div}S_{d}\rvert}{|Z|}=\frac{\lvert\operatorname{div}S\rvert}{|df|}$, where we omit composition with $\gamma$. To sum up, recalling that $\dot{\gamma}=\frac{(-\partial_{y}f,\partial_{x}f)}{|df|}$, we get
$\int_{0}^{L}\alpha_{d}\frac{(-\partial_{y}f-|\partial_{x}f|)^{+}}{|df|}\circ\gamma$ $\leq\Lambda\max\alpha_{d}+\frac{1}{\Lambda-2}\int_{0}^{\infty}\int_{\{\alpha_{d}>s\}}\frac{(-\partial_{y}f-|\partial_{x}f|)^{-}}{|df|}\circ\gamma(t)\,dt\,ds$
$\leq\Lambda\int_{0}^{L}\frac{\lvert\operatorname{div}S\rvert}{|df|}\circ\gamma+\frac{1}{\Lambda-2}\int_{0}^{L}\alpha_{d}\frac{(-\partial_{y}f-|\partial_{x}f|)^{-}}{|df|}\circ\gamma.$
@@
By approximation, it is enough to show the claim assuming that $\chi\in C^{0}_{c}(\mathbb{R}^{2})$ and (given such $\chi$) that $S,T\in C^{\infty}_{c}(\mathbb{R}^{2},\mathbb{R}^{2})$. Indeed, if $S_{\varepsilon},T_{\varepsilon}\rightharpoonup S,T$ as measures, with $|S_{\varepsilon}|,|T_{\varepsilon}|\rightharpoonup|S|,|T|$ and $\lvert\operatorname{div}S_{\varepsilon}\rvert,\lvert\operatorname{div}T_{\varepsilon}\rvert\rightharpoonup\lvert\operatorname{div}S\rvert,\lvert\operatorname{div}T\rvert$ tightly, then by Reshetnyak’s continuity principle we also have
@@
and similarly (since $|(S_{\varepsilon},T_{\varepsilon})|\rightharpoonup|(S,T)|$) also
@@
As in the previous proofs, we can also suppose that
$S=\alpha Z,\quad T=\beta W,$
for two smooth, complete, divergence-free vector fields $Z,W$ satisfying $Z^{x},W^{y}>0$ (for nonnegative coefficients $\alpha,\beta\in C^{\infty}_{c}(\mathbb{R}^{2})$). Also, we can assume that maximal integral curves of $Z$ are graphs over the (entire) horizontal axis, while those of $W$ are graphs over the vertical axis.
As in the previous subsection, we define the bands $\mathcal{U}_{j}^{x}:=[j,j+1]\times\mathbb{R}$ and $\mathcal{U}_{k}^{y}:=\mathbb{R}\times[k,k+1]$, and we let the sets $G_{j}^{x}$, $G_{k}^{y}$, and $G_{jk}=G_{j}^{x}\cap G_{k}^{y}$ be as in Definition 3.8.
Let us split again $S=S_{f}+S_{d}$ (see Eq. 17 and Eq. 19) on each band $\mathcal{U}_{j}^{x}$ and define $S_{f}^{P},S_{f}^{N},S_{d}^{P},S_{d}^{N}$ as in Eq. 11 (for instance, $S_{f}^{P}:=(S_{f}^{x}-|S_{f}^{y}|)^{+}$), as well as the analogous objects for $T$. Proposition 3.10 and Proposition 3.12 show that
$\int_{\mathcal{U}_{j}^{x}\setminus G_{j}^{x}}S_{f}^{P}\leq\int_{\mathcal{U}_{j}^{x}}S_{f}^{N}$
and
$\int_{\mathcal{U}_{j}^{x}}S_{d}^{P}\leq\Lambda\int_{\mathcal{U}_{j}^{x}}\lvert\operatorname{div}S\rvert+\frac{1}{\Lambda-2}\int_{\mathcal{U}_{j}^{x}}S_{d}^{N}.$
Setting $G^{x}:=\bigcup_{j\in\mathbb{Z}}G_{j}^{x}$ and summing the previous bounds over $j$, we then get
$\int_{\mathbb{R}^{2}\setminus G^{x}}S_{f}^{P}+\int_{\mathbb{R}^{2}}S_{d}^{P}\leq\Lambda\int_{\mathbb{R}^{2}}\lvert\operatorname{div}S\rvert+\int_{\mathbb{R}^{2}}S_{f}^{N}+\frac{1}{\Lambda-2}\int_{\mathbb{R}^{2}}S_{d}^{N}.$
To continue, recall that $S_{f}$ and $S_{d}$ are nonnegative multiples of $Z$. This implies that $S^{P}=S_{f}^{P}+S_{d}^{P}$ and $S^{N}=S_{f}^{N}+S_{d}^{N}$. Hence, fixing any $\Lambda\geq 3$ we get
(20) $\int_{\mathbb{R}^{2}\setminus G^{x}}S_{f}^{P}+\int_{\mathbb{R}^{2}}S_{d}^{P}\leq\Lambda\int_{\mathbb{R}^{2}}\lvert\operatorname{div}S\rvert+\int_{\mathbb{R}^{2}}S^{N}.$
Also, we can bound

### v2

@@ 3.3. Complementary inequalities
In the same setting as in the previous subsection, the vector field $S_{f}=\alpha_{f}Z$ admits a useful estimate on the complement $\partial_{L}\mathcal{R}^{x}\setminus G^{x}$, as we now show.
@@
$\int_{\mathcal{R}^{x}\setminus G^{x}}(S_{f}^{x}-|S_{f}^{y}|)^{+}\leq\int_{\mathcal{R}^{x}}(S_{f}^{x}-|S_{f}^{y}|)^{-},$
as well as the analogous one for $T_{f}$ on $\mathcal{R}^{y}$ (with $x$ and $y$ interchanged).
This fact will be a direct consequence of the next elementary lemmas. The first one is a version of the so-called rising sun lemma.
@@
Assume now $\xi(L)>\xi(0)$. The image $\xi(E)$ includes $\xi(0),\xi(L)]$ since, for any $\xi(0)\leq\lambda\leq\xi(L)$, we have $\max\xi^{-1}(\lambda)\in E$. Since $E\subseteq\{\dot{\xi}\geq 0\}$ up to negligible sets, the area formula gives
@@
The second lemma will be invoked several times and is thus stated in a generality that covers all the cases in which it will be used.
Lemma 3.12.
There exist two universal constants $\Lambda,C^{\prime}\geq 1$ such that the following hold, for a Lipschitz curve $\gamma:[0,L]\to\mathcal{U}_{j}^{x}=[j,j+1]\times\mathbb{R}$ with unit speed:
  • (i)
    if $\gamma^{x}(0)\geq\gamma^{x}(L)$ then
    $\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt\leq\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt;$
  • (ii)
    in general, without the previous assumption, we have
    $\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt\leq\Lambda+C^{\prime}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt.$
    If $\dot{\gamma}^{x}\geq 0$, then $C^{\prime}$ can be taken arbitrarily small (by increasing $\Lambda$).
Analogous statements hold for a stripe $\mathcal{U}_{k}^{y}$, interchanging the roles of $x,y$.
Proof.
The claim holds trivially in case (i), since
$\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)\,dt\leq\int_{0}^{L}\dot{\gamma}^{x}\,dt\leq 0$
and the left-hand side is the difference of the integrals of positive and negative parts. To check (ii), we can assume that $L\geq\Lambda$, since otherwise we can simply bound
$\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt\leq L\leq\Lambda.$
Let us fix $\theta_{0}\in(\frac{\pi}{4},\frac{\pi}{2})$ and $\alpha,\beta>0$ such that $(\cos\theta-|\sin\theta|)^{-}\geq\alpha$ for all $\theta\in[\theta_{0},\pi]$ and $\cos\theta\geq\beta$ for all $\theta\in[0,\theta_{0}]$. Letting $E:=\{t\,:\,\dot{\gamma}^{x}(t)\leq\cos\theta_{0}\}$ we have
$\alpha\cdot\mathcal{L}^{1}(E)\leq\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt,$
and hence
$\int_{[0,L]\setminus E}\dot{\gamma}^{x}\,dt\geq\beta\cdot\mathcal{L}^{1}([0,L]\setminus E)\geq\beta L-\frac{\beta}{\alpha}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt,$
while trivially
$\int_{[0,L]\setminus E}\dot{\gamma}^{x}\,dt\leq\int_{0}^{L}\dot{\gamma}^{x}\,dt+\mathcal{L}^{1}(E)\leq 1+\frac{1}{\alpha}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt,$
giving
$\frac{\beta+1}{\alpha}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt\geq\beta L-1.$
Once we fix any $\Lambda\geq\frac{2}{\beta}$, we have $\beta\Lambda-1\geq\frac{\beta\Lambda}{2}\geq\frac{\beta L}{2}$ and thus the bound
$\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt\leq L\leq\frac{2(\beta+1)}{\alpha\beta}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\,dt,$
giving again the claim.
If $\dot{\gamma}^{x}\geq 0$ we can argue instead as follows: we have $\dot{\gamma}^{x}+|\dot{\gamma}^{y}|\geq 1$ and thus
$\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}\geq\int_{0}^{L}(|\dot{\gamma}^{y}|-\dot{\gamma}^{x})\geq\int_{0}^{L}(1-2\dot{\gamma}^{x})\geq\Lambda-2,$
which gives
$\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\leq\int_{0}^{L}\dot{\gamma}^{x}\leq 1\leq\frac{1}{\Lambda-2}\int_{0}^{L}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}.\qed$
@@
As above, we write $S_{f}=\nabla^{\perp}\varphi$. Given a level set of $\varphi$, we can parametrize it with an integral curve $\gamma:[0,L]\to\mathcal{R}^{x}$ of $Z/|Z|$, with $\gamma^{x}(0)=j$ and $\gamma^{x}(L)=j+1$. Note that $\gamma$ has unit speed and that
@@
We apply 3.11 with $\xi(t):=\int_{0}^{t}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)$. With $E$ as in the statement of the lemma, if $s\in E$ then for all $t>s$ we have
@@
Integrating over all level sets of $\varphi$ and using the coarea formula, we get
$\int_{\mathcal{R}^{x}\setminus G_{j}^{x}}(-\partial_{y}\varphi-|\partial_{x}\varphi|)^{+}\leq\int_{\mathcal{R}^{x}}(-\partial_{y}\varphi-|\partial_{x}\varphi|)^{-}.$
@@
Recall that $0<\alpha_{f}\leq\alpha$, which gives $0\leq\alpha_{d}<\alpha$.
@@
For the same constants $\Lambda,C^{\prime}$ as above, we have
$\int_{\mathcal{R}^{x}}(S_{d}^{x}-|S_{d}^{y}|)^{+}\leq\Lambda\int_{\mathcal{R}^{x}}\lvert\operatorname{div}S\rvert+C^{\prime}\int_{\mathcal{U}_{j}^{x}}(S_{d}^{x}-|S_{d}^{y}|)^{-}.$
If $S^{x}\geq 0$ then we can take $C^{\prime}=1$.
@@
Writing $Z=\nabla^{\perp}f$, consider a level set $\gamma:[0,L]\to\mathcal{R}^{x}$ of $f$, parametrized by arclength. We have
@@
where we write $\alpha_{d}$ in place of $\alpha_{d}\circ\gamma$ for simplicity. Note that, calling $N_{s}$ the cardinality of $\{\alpha_{d}=s\}$, the set $\{\alpha_{d}>s\}$ has at most $N_{s}$ connected components, since $\alpha_{d}$ takes all values in $(0,\max\alpha_{d})$. Hence, applying part (ii) of 3.12 to $\gamma$ restricted to each connected component, we get
$\int_{\{\alpha_{d}>s\}}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt\leq\Lambda N_{s}+C^{\prime}\int_{\{\alpha_{d}>s\}}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\,dt.$
Integrating in $s$ and using the area formula, we deduce that
$\int_{0}^{L}\alpha_{d}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{+}\leq\Lambda\int_{0}^{L}|\dot{\alpha}_{d}|+C^{\prime}\int_{0}^{L}\alpha_{d}(\dot{\gamma}^{x}-|\dot{\gamma}^{y}|)^{-}.$
If $S^{x}\geq 0$ then $C^{\prime}$ can be taken as small as desired, as $\dot{\gamma}^{x}\geq 0$ in this case; in particular, we can take $C^{\prime}=1$ (with $\Lambda:=3$).
Since $\dot{\gamma}=\frac{Z}{|Z|}$ and $\operatorname{div}Z=0$, we have $|\dot{\alpha}_{d}|=\frac{\lvert\operatorname{div}S_{d}\rvert}{|Z|}=\frac{\lvert\operatorname{div}S\rvert}{|df|}$, where we omit composition with $\gamma$. To sum up, recalling that $\dot{\gamma}=\frac{(-\partial_{y}f,\partial_{x}f)}{|df|}$, we get
$\int_{0}^{L}\alpha_{d}\frac{(-\partial_{y}f-|\partial_{x}f|)^{+}}{|df|}\circ\gamma$ $\leq\Lambda\int_{0}^{L}\frac{\lvert\operatorname{div}S\rvert}{|df|}\circ\gamma+C^{\prime}\int_{0}^{L}\alpha_{d}\frac{(-\partial_{y}f-|\partial_{x}f|)^{-}}{|df|}\circ\gamma.$
@@
Remark 3.14.
Note that the notion of good region $\mathcal{R}^{x}$ for $S$ can be relaxed slightly: we can rather assume that there exists a smooth unit vector field $U$ on $\mathcal{R}^{x}$ whose integral curves foliate $\mathcal{R}^{x}$ and that (pointwise) $S$ is a nonnegative multiple of $U$. The previous proofs still apply: we can repeat them for $S+\delta U$ (with $\delta>0$) and then let $\delta\to 0$.
@@
By approximation, it is enough to show the claim assuming that $\chi\in C^{0}_{c}(\mathbb{R}^{2})$ and (given such $\chi$) that $S,T\in C^{\infty}_{c}(\mathbb{R}^{2},\mathbb{R}^{2})$. Indeed, if $S_{\varepsilon},T_{\varepsilon}\rightharpoonup S,T$ as measures, with $|(S_{\varepsilon},T_{\varepsilon})|\rightharpoonup|(S,T)|$ and $\lvert\operatorname{div}S_{\varepsilon}\rvert,\lvert\operatorname{div}T_{\varepsilon}\rvert\rightharpoonup\lvert\operatorname{div}S\rvert,\lvert\operatorname{div}T\rvert$ tightly, then by Reshetnyak’s continuity principle we also have
@@
and similarly
@@
While reference vector fields $Z,W$ as above can no longer be constructed on all the plane, we first claim that we can reduce ourselves to the following situation, for a suitable finite set $F\subset\mathbb{R}^{2}$:
  • (i)
    $S=\eta U$ and $T=\zeta V$, where $\eta,\zeta\in C^{\infty}_{c}(\mathbb{R}^{2}\setminus F)$;
  • (ii)
    $U,V:\mathbb{R}^{2}\setminus F\to\mathbb{S}^{1}$ are smooth;
  • (iii)
    $U=\partial_{x}$ and $V=\partial_{y}$ outside of a large disk;
  • (iv)
    for each $p\in F$, each of $U,V$ has the following property: either it extends smoothly across $p$ or it has the form $\pm\frac{i(z-p)}{|z-p|}$ near $p$, in complex notation.
The last condition implies that all but finitely many integral curves of $U$ and $V$ are defined at all times $t\in\mathbb{R}$ (and are disjoint from $F$).
To see this claim, let $\chi_{\varepsilon}:\mathbb{R}^{2}\to\mathbb{R}$ be a nonnegative smooth function such that $\chi_{\varepsilon}=0$ on $\overline{B}_{\varepsilon}(0)$ and $\chi_{\varepsilon}=1$ on $\mathbb{R}^{2}\setminus B_{2\varepsilon}$, with $|d\chi_{\varepsilon}|\leq 2/\varepsilon$. Letting $S_{\varepsilon}:=\chi_{\varepsilon}(S)S$ and $T_{\varepsilon}:=\chi_{\varepsilon}(T)T$, it is immediate to check that $S_{\varepsilon},T_{\varepsilon},\operatorname{div}S_{\varepsilon},\operatorname{div}T_{\varepsilon}$ converge to $S,T,\operatorname{div}S,\operatorname{div}T$ in $L^{1}(\mathbb{R}^{2})$, as $\varepsilon\to 0$, so that we reduce ourselves to show the bound for $S_{\varepsilon},T_{\varepsilon}$. Taking now $\varepsilon>0$ to be a regular value for both $|S|$ and $|T|$, we define
$\Omega_{\varepsilon}:=\{|S|>\varepsilon\},$
which is a bounded smooth domain, and we make the following simple observation.
Lemma 3.15.
The vector field $S/|S|$, defined on $\Omega_{\varepsilon}$, can be extended to a smooth vector field $U_{\varepsilon}:\mathbb{R}^{2}\setminus F_{\varepsilon}\to\mathbb{S}^{1}$ such that we can write $S_{\varepsilon}=\eta_{\varepsilon}U_{\varepsilon}$, with $F_{\varepsilon},\eta_{\varepsilon},U_{\varepsilon}$ satisfying all the requirements in the claim above.
Proof.
Taking $R_{\varepsilon}>0$ such that $\overline{\Omega_{\varepsilon}}$ is included in an open disk $B_{R_{\varepsilon}}(0)$, we let $\Omega_{\varepsilon}^{\prime}:=\Omega_{\varepsilon}\cup(\mathbb{R}^{2}\setminus\overline{B}_{R_{\varepsilon}}(0))$ and initially define $U_{\varepsilon}$ on its closure, by letting $U_{\varepsilon}:=S/|S|$ on $\overline{\Omega_{\varepsilon}}$, while $U_{\varepsilon}:=\partial_{x}$ on $\mathbb{R}^{2}\setminus B_{R_{\varepsilon}}(0)$.
We can then extend it to a vector field $U_{\varepsilon}:\Omega_{\varepsilon}^{\prime\prime}\to\mathbb{S}^{1}$ (smooth up to the boundary), for a connected $\Omega_{\varepsilon}^{\prime\prime}\supseteq\Omega_{\varepsilon}^{\prime}$ (with smooth boundary): for instance, we can find finitely many disjoint injective curves $\gamma_{j}\subset\mathbb{R}^{2}\setminus\Omega_{\varepsilon}^{\prime}$ (meeting $\partial\Omega_{\varepsilon}^{\prime}$ transversely at the endpoints) such that $\Omega_{\varepsilon}^{\prime}\cup\bigcup_{j}\gamma_{j}$ becomes connected; each $\gamma_{j}$ admits a neighborhood $\Omega_{j}$ diffeomorphic to a $\delta_{j}$-neighborhood $\Delta_{j}$ of the unit segment $0,1]\times\{0\}$, through a diffeomorphism $\Psi$ which maps $\Omega_{\varepsilon}\cap\Omega_{j}$ to $\{(x,y)\in\Delta_{j}\,:\,x\not\in[0,1]\}$. Viewing $\mathbb{S}^{1}$ as unit complex numbers, we can write $U_{\varepsilon}\circ\Psi^{-1}$ as $e^{i\varphi_{j}}$ on the latter set (as it has two simply connected components), extend $\varphi_{j}$ smoothly on all of $\Delta_{j}$, and then define $U_{\varepsilon}:=e^{i\varphi_{j}}\circ\Psi$ on all of $\Omega_{j}$. It then suffices to replace $\Omega_{\varepsilon}^{\prime}\cup\bigcup_{j}\Omega_{j}$ with a subset $\Omega_{\varepsilon}^{\prime\prime}$ with smooth boundary including $\Omega_{\varepsilon}^{\prime}\cup\bigcup_{j}\gamma_{j}$.
Finally, since $\Omega_{\varepsilon}^{\prime\prime}$ is connected, its complement is diffeomorphic to the union of finitely many disjoint closed disks. Identifying each of them with the closed unit disk $\overline{D}$, we can extend $U_{\varepsilon}$ from the boundary $\partial D$ to $\overline{D}\setminus\{z_{1},\dots,z_{|k|}\}$ (for arbitrarily chosen distinct points $z_{1},\dots,z_{|k|}\in D$), where $k\in\mathbb{Z}$ is the degree of $U_{\varepsilon}$ on $\partial D$, in such a way that the extension has degree $\pm 1$ around $z_{\ell}$, for each $\ell=1,\dots,|k|$ (namely, $1$ if $k>0$ and 1-1 if $k<0$); in fact, we can choose $U_{\varepsilon}$ to have the form $\pm\frac{i(z-z_{\ell})}{|z-z_{\ell}|}$ near each $z_{\ell}$. ∎
We repeat the same for $T_{\varepsilon}$, obtaining $T_{\varepsilon}=\zeta_{\varepsilon}V_{\varepsilon}$ for a suitable $V_{\varepsilon}:\mathbb{R}^{2}\setminus F_{\varepsilon}^{\prime}\to\mathbb{S}^{1}$. In the sequel, we let $F:=F_{\varepsilon}\cup F_{\varepsilon}^{\prime}$ and we simply write $S,T,U,V$ in place of $S_{\varepsilon},T_{\varepsilon},U_{\varepsilon},V_{\varepsilon}$.
Up to a slight rotation in the values of $S$ (and $U$), we can also assume that $(0,\pm 1)$ are regular values for $U$ on the line $\{j\}\times\mathbb{R}$, for each $j\in\mathbb{Z}$. Thus, up to adding finitely many points to $F$, we can assume that $U(x,y)\neq(0,\pm 1)$ whenever $x\in\mathbb{Z}$ and $(x,y)\not\in F$, and similarly that $V(x,y)\neq(\pm 1,0)$ whenever $y\in\mathbb{Z}$ and $(x,y)\not\in F$.
Next, let us consider again the vertical and horizontal stripes introduced previously, namely $\mathcal{U}_{j}^{x}=[j,j+1]\times\mathbb{R}$ and $\mathcal{U}_{k}^{y}=\mathbb{R}\times[k,k+1]$. By construction, up to a negligible set, each stripe $\mathcal{U}_{j}^{x}$ is a union of (pieces of) integral curves of $U$ given by one of the following four kinds:
  • (i)
    curves joining the left side to the right side, in this order;
  • (ii)
    curves joining the right side to the left one, or starting from one side and ending on the same one;
  • (iii)
    injective curves of infinite length;
  • (iv)
    loops;
also, each of these meets $\partial\mathcal{U}_{j}^{x}$ transversely.
The union of curves of type (i) forms an open set, determined by a set of good initial conditions $\{j\}\times\bigcup_{\ell}I_{j,\ell}$ on the left side, for disjoint open intervals $I_{j,\ell}$. We let $\mathcal{R}_{j,\ell}^{x}\subseteq\mathcal{U}_{j}^{x}$ be the union of those (pieces of) integral curves with initial condition in $\{j\}\times I_{j,\ell}$, which is a good region for $S$ according to Definition 3.7 (see 3.14).
Repeating the same on horizontal stripes, finding regions $\mathcal{R}_{k,m}^{y}\subseteq\mathcal{U}_{k}^{y}$ where integral curves of $V$ travel from the bottom side to the top one, we can define $G_{j,\ell}^{x}\subseteq\mathcal{R}_{j,\ell}^{x}$ and $G_{k,m}^{y}\subseteq\mathcal{R}_{k,m}^{y}$ as in Definition 3.8, relative to these good regions. Thus, by Eq. 15 we have
[…]
Letting $G^{x}:=\bigcup_{j,\ell}G_{j,\ell}^{x}$, $G^{y}:=\bigcup_{k,m}G_{k,m}^{y}$, and $G:=G^{x}\cap G^{y}$, we deduce that
[…]
where $S_{f},T_{f}$ are defined piecewise on each good region for $S$ and $T$, respectively, as in Eq. 16. Note that $S_{f}^{P}T_{f}^{P}\leq C\det(S_{f},T_{f})$ on $G$ by the elementary bound
$(v^{x}-|v^{y}|)(w^{y}-|w^{x}|)\leq C\det(v,w)$
for vectors $v\in\overline{C^{x}}$ and $w\in\overline{C^{y}}$ (note that $S_{f}\in\overline{C^{x}}$ on $G$, except possibly on the negligible set $\mathbb{Z}\times\mathbb{R}$, and similarly $T_{f}\in\overline{C^{y}}$). In turn, this bound follows from the fact that, assuming $|v|=|w|=1$, $\det(v,w)$ is comparable with $\min\{|v-w|,|v+w|\}$, while $v^{x}-|v^{y}|$ is comparable with $\operatorname{dist}(v,\partial C^{x})$, and similarly $w^{y}-|w^{x}|$ is comparable with $\operatorname{dist}(w,\partial C^{y})$. From this remark and Cauchy–Schwarz it follows that
$\int_{G}\chi\sqrt{S_{f}^{P}T_{f}^{P}}\leq C\Big[\int_{\mathbb{R}^{2}}(S^{x}+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(T^{y}+\lvert\operatorname{div}T\rvert)\Big]^{1/2}.$
Recalling that $S_{f},S_{d}$ are nonnegative multiples of $U$, we get $S^{P}=S_{f}^{P}+S_{d}^{P}$ and $S^{N}=S_{f}^{N}+S_{d}^{N}$, and hence